OpenStudy (anonymous):

Find the derivatives of the function f(t)=cosh2(4t)−sinh2(4t). Simplify your answer completely.

OpenStudy (zarkon):

is it \[\cosh^2(4t)-\sinh^2(4t)\]

OpenStudy (anonymous):

yep. sorry.

OpenStudy (zarkon):

\[\cosh^2(4t)-\sinh^2(4t)=1\]

OpenStudy (anonymous):

oh, haha. I guess that makes sense.

OpenStudy (zarkon):

and the derivative of a constant is zero

OpenStudy (anonymous):

It said 1 is wrong. Would it be 2sinh(4t)+2cosh(4t)?

OpenStudy (zarkon):

I didn't say 1 was the answer

OpenStudy (zarkon):

\[\frac{d}{dt}\left[\cosh^2(4t)-\sinh^2(4t)\right]=\frac{d}{dt}[1]=0\]

OpenStudy (anonymous):

Ok, so 2cosh(4t)*4sinh(4t)-2sinh(4t)*4cosh(4t)?

OpenStudy (zarkon):

which is zero

OpenStudy (anonymous):

Ok.